Pixel:

The equation y2−4y−x+9=0 represents a parabola. Rewrite the equation of this parabola in standard form.

Pixel:

1 attachment
dude:

Not +9

Pixel:

is it y-2

dude:

Ye that part is right

Pixel:

-9

dude:

No You just need to complete the square here \(y^2-4y+~? = x-9+~?\) \(ax^2+bx+c\) -- Left side of equation Basically use \((\dfrac{b}{2})^2\) to find what to add [= 4] Use \((\dfrac{b}{2})\) to find the root/factor [Which was -2] \(y^2-4y+4 = x-9+4\) Now factoring \((y-2)^2= x -5\) And now just get the -5 to the other side of the eq => then answer

Pixel:

+5

dude:

Ye